OpenStudy (anonymous):

What is the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6?

rishavraj (rishavraj):

if two lines are parallel it means their slopes are equal(same) the given line is 3x - 2y = -6 slope of this line = 3/2 now proceed

OpenStudy (anonymous):

so it would be 3/2-2y=-6

OpenStudy (anonymous):

@rishavraj

rishavraj (rishavraj):

nah nah its not tht see now u know slope of the line....it is m = 3/2 \[\frac{ y_2 - y_1 }{ x_2 - x_1 } = m \]